OpenStudy (shaik0124):

n a division problem, the divisor is twenty times the quotient and five times the remainder. If remainder is 16, the number will be: Option 1 : 3360 Option 2 : 336 Option 3 : 1616 Option 4 : 20516

ganeshie8 (ganeshie8):

\[n = d*q + r\] use the given relationships and see if you can figure out

OpenStudy (shaik0124):

given is d= 20q+5r and r=16 d=20q+16*5 d=20q+80 then?

ganeshie8 (ganeshie8):

Not exactly... we're given \(r=16\)

ganeshie8 (ganeshie8):

` the divisor is twenty times the quotient and five times the remainder` that translates to \[d = 20q = 5r\] yes ?

OpenStudy (ribhu):

336

OpenStudy (shaik0124):

n=80*4+16 n=320+16 336

ganeshie8 (ganeshie8):

looks good!

OpenStudy (ribhu):

thanks @ganeshie8